# Anomaly Detection in Large-Scale Cloud Systems: An Industry Case and Dataset

### v1

@@ I Introduction
  • Demonstrating predictive models for anomaly detection in cloud environments.
@@ II-B Anomaly Detection Methods
In our prior work [9, 10], we proposed an anomaly detection architecture based on a GRU-based autoencoder with a likelihood function for anomaly detection in multi-dimensional cloud telemetry. Our initial results showed that this model could detect anomalies up to 20 minutes earlier than previous monitoring systems and significantly reduced false alerts. The detector’s performance was validated against both publicly available benchmark datasets and real-world data from the IBM Cloud Platform. Since 2021, the IBM Cloud Platform has undergone significant changes, including migration to containerized environments and the implementation of new technologies, leading to data with new characteristics that require tailored anomaly detection methods to address evolving challenges effectively. In this paper, we aim to highlight the challenges raised by the IBM Cloud Platform’s evolved system, particularly in working with the new data it generates. Additionally, we examine the difficulties in identifying ground truth labels and constructing automatic anomaly detection techniques on these complex datasets.
@@ III-A Software System Under Study
The Console production software system is deployed across seven data centers worldwide, though not all instances are active at the same time; they are rotated based on operational requirements. Additional Console deployments are spread globally for testing and staging environments. The Console uses a microservices architecture, with each microservice generating millions of daily logs and telemetry records (a common challenge with large scale Cloud systems [42, 43, 44, 45]). These records offer valuable insights into the health and performance of the IBM Cloud, providing critical data for monitoring and optimization.
@@ IV-B2 Evaluation Metrics
For a detailed explanation of the NAB score, refer to [21, 17]. In brief, this metric rewards early detection (TP) within an anomaly window and penalizes false positives and false negatives. Any detection within an anomaly window is considered a TP, with its score determined by a scaled sigmoid function. A detection at the start of the anomaly window is assigned a score of 1, while detections near the end receive lower scores, closer to 0. Any subsequent detections within the same window are ignored for scoring purposes, meaning that all detections within one window count as a single TP. If the model misses an entire anomaly window, it results in one FN. FP and TN, on the other hand, are evaluated point-by-point. TNs do not affect NAB score, but each FP lowers the score.
The NAB score can be computed using different cost profiles; we use two shown in Table I: “Standard” and “Reward Low FN”. The “Standard” profile strikes a balance between TP, FN, and FP, applying a modest penalty on FP to prevent the model from favoring either precision or recall too heavily. In contrast, the “Reward Low FN” profile applies a harsher penalty on FN, which is useful in scenarios like ours where detecting each anomaly is crucial due to the importance of anomalies present in the dataset. This higher penalty on false negatives encourages the model to prioritize finding the anomalies, even at the risk of increasing FP count.
TABLE I: NAB score cost profiles as per [17].
@@ IV-C Results
TABLE II: Performance of the anomaly detectors. AE denotes autoencoder.
@@ IV-C1 Quantitative Results
Table II presents the performance of the models in detecting anomalies.
The GRU model showed a slight improvement over the ANN model. The NAB score increased from 4.61 to 6.06 for the “Standard Profile” and from 4.61 to 13.60 for the “Low FB Profile.” This result aligns with expectations, as the GRU architecture is inherently more powerful than the ANN.
@@ V Insights and Challenges
The size of the system and the corresponding large feature set (on the order of $10^{5}$) make model training computationally expensive. This, in turn, complicates model hyperparameter tuning due to the high resource demands. One possible solution is to reduce the dimensionality, either by selecting salient features based on expert knowledge or by employing automatic dimensionality reduction techniques.
@@ VI-A External Validity
Software engineering studies are often challenged by the variability of real-world environments and the problem of generalization cannot be fully resolved [69]. In this work, we consider our software and its associated telemetry dataset as a critical case [^7] In case study research, a critical case refers to a particularly significant or decisive instance chosen for its potential to offer deep insights into a specific phenomenon [68]., which could assist the community in building anomaly detection models and testing them on the data from a real-world LCS. The same empirical evaluation and analysis methods can be applied to other software products, provided that well-designed and controlled experiments are conducted.
@@ B-A Aggregated telemetry
  • statusCode: $30$ distinct HTTP response status codes [74] (e.g., 200 or 500) and $1$ non-HTTP response status code (namely, -1). The count of columns for different groups of HTTP status codes in the pivoted version of the dataset is given in Table III.
@@
TABLE III: Number of columns for different groups of HTTP status codes in the pivoted version of the dataset.

### v2

@@ I Introduction
  • Demonstrating predictive models for anomaly detection in cloud environments. The reproducibility package is accessible on GitHub [21] and Zenodo[22].
@@ II-B Anomaly Detection Methods
In our prior work [9, 10], we proposed an anomaly detection architecture based on a GRU-based autoencoder with a likelihood function for anomaly detection in multi-dimensional cloud telemetry. Our initial results showed that this model could detect anomalies up to 20 minutes earlier than previous monitoring systems and significantly reduced false alerts. The detector’s performance was validated against both publicly available benchmark datasets and real-world data from the IBM Cloud Platform. That work focused on an earlier version of the software under study, which has significantly evolved over the past few years, becoming more sophisticated yet increasingly complex (in line with the Laws of Software Evolution [43]). Since 2021, the IBM Cloud Platform has undergone significant changes, including migration to containerized environments and the implementation of new technologies, leading to data with new characteristics that require tailored anomaly detection methods to address evolving challenges effectively.
In this paper, we aim to highlight the challenges raised by the IBM Cloud Platform’s evolved system, particularly in working with the new data it generates. Unlike the previous studies [9, 10], which did not share the data or models, this work provides both, enabling the community and industry to adopt the dataset as a benchmark for detection of anomalies in operation of large-scale anomaly cloud software. Additionally, we examine the difficulties in identifying ground truth labels and constructing automatic anomaly detection techniques on these complex datasets.
@@ III-A Software System Under Study
The Console production software system is deployed across seven data centers worldwide, though not all instances are active at the same time; they are rotated based on operational requirements. Additional Console deployments are spread globally for testing and staging environments. The Console uses a microservices architecture, with each microservice generating millions of daily logs and telemetry records (a common challenge with large scale Cloud systems [45, 46, 47, 48]). These records offer valuable insights into the health and performance of the IBM Cloud, providing critical data for monitoring and optimization. The sheer volume and variability of logs and telemetry, combined with the dynamic nature of containerized environments — characterized by transient workloads, volatile resource usage, and continuous scaling — necessitate advanced, context-aware anomaly detection techniques. Traditional static or threshold-based methods fall short in addressing the complexity and scale of these systems. Our anomaly detection approach aims to address these challenges by leveraging innovative techniques designed for dynamic environments, contributing to more effective monitoring and actionable insights.
@@ III-E Comparison with Existing Datasets
TABLE I: Comparison of Datasets for Anomaly Detection in Cloud Systems.
Dataset File Column Avg. Row Row
Name Count Count Count Range
NAB $57$ $1$ $6303$ 112711271127112722 6952269522\,69522 695
MS Cloud Monitoring $67$ $1$ $3757$ 17617617617620 1602016020\,16020 160
Exathlon $93$ $2283$ 25 2002520025\,20025 200 25 2002520025\,20025 200
Ours $1$ 117 449117449117\,449117 449 39 3653936539\,36539 365 39 3653936539\,36539 365
@@ IV-B2 Evaluation Metrics
For a detailed explanation of the NAB score, refer to [23, 17]. In brief, this metric rewards early detection (TP) within an anomaly window and penalizes false positives and false negatives. Any detection within an anomaly window is considered a TP, with its score determined by a scaled sigmoid function. A detection at the start of the anomaly window is assigned a score of 1, while detections near the end receive lower scores, closer to 0. Any subsequent detections within the same window are ignored for scoring purposes, meaning that all detections within one window count as a single TP. If the model misses an entire anomaly window, it results in one FN. FP and TN, on the other hand, are evaluated point-by-point. TNs do not affect the NAB score, but each FP lowers the score.
The NAB score can be computed using different cost profiles; we use two shown in Table II: “Standard” and “Reward Low FN”. The “Standard” profile strikes a balance between TP, FN, and FP, applying a modest penalty on FP to prevent the model from favoring either precision or recall too heavily. In contrast, the “Reward Low FN” profile applies a harsher penalty on FN, which is useful in scenarios like ours where detecting each anomaly is crucial due to the importance of anomalies present in the dataset. This higher penalty on false negatives encourages the model to prioritize finding the anomalies, even at the risk of increasing FP count.
TABLE II: NAB score cost profiles as per [17].
@@ IV-C Results
TABLE III: Performance of the anomaly detectors. AE denotes autoencoder.
@@ IV-C1 Quantitative Results
Table III presents the performance of the models in detecting anomalies.
The GRU model showed a slight improvement over the ANN model. The NAB score increased from 4.61 to 6.06 for the “Standard Profile” and from 4.61 to 13.60 for the “Low FN Profile.” However, the marginal improvements observed with the GRU model were somewhat unexpected, suggesting potential for further enhancement through hyperparameter tuning. Techniques such as advanced optimization strategies or refinements to the model architecture could better harness the GRU’s capabilities and address current limitations in the experimental setup.
@@ V Insights and Challenges
The size of the system and the corresponding large feature set (on the order of $10^{5}$) make model training computationally expensive. This, in turn, complicates model hyperparameter tuning due to the high resource demands. For example, during our experiments on the complete dataset, hyperparameter tuning for the trained models demanded significant computational resources, often resulting in prolonged training times and system bottlenecks, such as memory exhaustion or CPU overload during the deployment of anomaly detection systems. One possible solution is to reduce the dimensionality, either by selecting salient features based on expert knowledge or by employing automatic dimensionality reduction techniques. Developing scalable algorithms that balance computational efficiency and model performance remains a critical area for future work.
@@ VI-A External Validity
Software engineering studies are often challenged by the variability of real-world environments, and the problem of generalization cannot be fully resolved [72]. In this work, we consider our software and its associated telemetry dataset as a critical case [^7] In case study research, a critical case refers to a particularly significant or decisive instance chosen for its potential to offer deep insights into a specific phenomenon [71]., which could assist the community in building anomaly detection models and testing them on the data from a real-world LCS. The same empirical evaluation and analysis methods can be applied to other software products, provided that well-designed and controlled experiments are conducted.
Beyond addressing the generalization problem, this dataset holds practical value for various stakeholders in cloud environments. For instance, if a client operates a complex cloud application (e.g., distributed and comprising multiple components), this dataset would remain relevant and of interest. Models trained at the provider level can help customers by identifying patterns or anomalies specific to distributed workloads, thereby improving the robustness and reliability of their operations. This highlights the practical value of the dataset beyond cloud providers, extending its applicability to end-users managing sophisticated cloud-based applications.
@@ B-A Aggregated telemetry
  • statusCode: $30$ distinct HTTP response status codes [77] (e.g., 200 or 500) and $1$ non-HTTP response status code (namely, -1). The count of columns for different groups of HTTP status codes in the pivoted version of the dataset is given in Table IV.
@@
TABLE IV: Number of columns for different groups of HTTP status codes in the pivoted version of the dataset.